OpenStudy (anonymous):

Evaluate the series 72 + 12 + 2 + . . . a.14.4 b.432 c.6 d.86.4

OpenStudy (blurbendy):

72 / 6 = 12 12/6 = 2 2/6 = 1/3 1/3 / 6, etc ,etc

OpenStudy (blurbendy):

I believe it will converge to 86.4

terenzreignz (terenzreignz):

This is an infinite geometric series, starting at 72 with a common ratio 1/6

OpenStudy (anonymous):

so d?

terenzreignz (terenzreignz):

Formula is \[\Large S = \frac{a}{1-r}\] where a is the first term and r is the common ratio...

terenzreignz (terenzreignz):

Yeah, D seems right :)